Lydia bought a stack of magazines to resell for profit. She sold half of them on Tuesday. She sold 10 on Wednesday. She sold half of what was left on Thursday. On Friday she sold the remaining 8 magazines. How many magazines did she begin with?

Answer by stanbon(75887)

Lydia bought a stack of magazines to resell for profit. She sold half of them on Tuesday. She sold 10 on Wednesday. She sold half of what was left on Thursday. On Friday she sold the remaining 8 magazines. How many magazines did she begin with?
-----------------------------
Let original number be "x".
Equation:
(1/2)[x - (1/2)x-10] = 8
-------------------
16 = (1/2)x-10
(1/2)x = 26
x = 52
-----
Checking:
(1/2)[52-26-10) = 8
16 = 16
==================
Cheers,
Stan H.
==================